OpenStudy (anonymous):

Help solving Find the critical value z a/2 that corresponds to a=0.09

OpenStudy (amoodarya):

I think your question is not clear !

OpenStudy (anonymous):

no its exactly how its written on my homework problem @amoodarya

OpenStudy (anonymous):



OpenStudy (anonymous):

so you want to find the Z value such that \(P(|Z|>z_{\alpha/2})=\alpha\); the equivalent one-tailed probability is just half of that: \(P(Z>z_{\alpha/2})=\alpha/2\) so you can look at the cdf \(P(Z<z_{\alpha/2})=1-\alpha/2\) and apply an inverse Z cdf (or equivalently look up the \(z_{\alpha/2}\) corresponding to an area of \(1-\alpha/2\) in your Z table)